Olaf is standing on a sheet of ice that covers the football stadium parking lot in Buffalo, New York; there is negligible friction between his feet and the ice. A friend throws Olaf a ball of mass 0.400 kg that is traveling horizontally at 11.8 m/s . Olaf's mass is 73.4 kg.

A. If Olaf catches the ball, with what speed vf do Olaf and the ball move afterward?

B. If the ball hits Olaf and bounces off his chest horizontally at 8.80 m/s in the opposite direction, what is his speed vf after the collision?
Answer #1

momentum conservation of olaf+ball

Pi = Pf

0.400*11.8 = (73.4+0.4)*V

=> V = vf = 0.063 m/s

B.

momentum conservation of olaf+ball

0.400*11.8 = 73.4*V - 0.400*8.8

=> V = 0.112